# PerfTracker: Online Performance Diagnosis for Large Model Training

### v3

@@ 5. Implementation
We built PerfTracker with 6K lines of Python code, and deployed PerfTracker as a service in our production GPU clusters. Deploying PerfTracker in production GPU clusters required us to address the following practical challenges.
@@
Optimizations of profiling data generation. PerfTracker relies on Torch profiler to generate a part of raw profiling data (§4). However, Torch profiler introduces overhead to LMT: (1) When the profiling window ends, it requires a long time to prepare and dump the data, and (2) after profiling, it leaves a lot of resources (e.g., hooks for CUDA functions added by CUPTI (cup, 2024)) in the LMT. They may reduce the performance of CUDA kernel execution even after profiling is finished.
@@
PerfTracker addresses the problems by optimizing the Torch profiler. First, the Torch profiler always transfers profiling data to Chrome tracing format, and then dumps the data using Kineto APIs (kin, 2024). Since Kineto supports dumping data with the same format, we remove the redundant and slow format transformation performed by Torch profiler, and directly apply Kineto to dump data. This reduces the time spent on data generation by 33%. Second, PerfTracker calls cuptiFinalize() to clear all remaining resources in LMT after profiling.
@@
Accessing hardware information within the user’s container. In production GPU clusters, an LMT typically runs inside containers; so does PerfTracker daemons (§4.1). However, GPU cluster providers usually restrict the permissions of user containers; therefore, PerfTracker cannot directly access hardware information. PerfTracker utilizes Kubernetes (kub, 2024)’s native “emptyDir” feature to share directories between containers, performing high-frequency hardware sampling through a privileged management container, and placing the data in the shared directory. Therefore, PerfTracker can obtain hardware profiling data within user containers without loosening permissions of user containers.
@@
Cooperating with other hardware monitors. In most production GPU clusters, coarse-grained hardware monitors are deployed on each physical host, using APIs such as DCGM and PCM to collect hardware metrics in real-time for health checks. However, some information (e.g., GPU metrics) can only be subscribed by one process at a time. PerfTracker’s data generator cooperates with the host’s monitoring system via signals (through a shared directory). As PerfTracker’s profiling typically lasts about 20 seconds each time, we do not encounter any conflicts caused by the cooperation.
6. Evaluation in Production
@@
PerfTracker has been deployed in our production GPU clusters with $O(10,000)$ GPUs for three months. During this period, there were in total dozens of LMT performance issues for which we failed to localize the root causes using state-of-the-art techniques described in §2.2 (see Table 2). PerfTracker successfully identified root causes of all of these performance issues, improving the training throughput of these LMTs (the largest one includes 3,072 GPUs) by $\sim$20to $\sim$100%.
@@
We present how we used PerfTracker to diagnose two LMTs with 2,560 and 3,072 GPUs respectively, to show:
@@
    PerfTracker can identify real-world performance issues of both hardware and code-level problems. (§6.1-6.2)
@@
    PerfTracker makes diagnosis in high efficiency, and does not introduce overhead to the routine training of LMT. (§6.3)
@@
    The comparison with the state of the arts. (§6.4)
@@
Table 2. Issues identified by PerfTracker (only includes performance issues not identified by our existing systems)
Category Root cause Proportion of LMT cases
Hardware issues GPU 8.3%
Network 16.7%
Misconfigurations PyTorch 8.3%
Communication 16.7%
Dataloader 8.3%
Low-efficiency code of users 41.7%
6.1. Case 1: Hardware Issues
@@
A text-to-picture LMT job of 2,560 Nvidia H800 GPUs is expected to take 5 seconds per iteration, but is training at 9 seconds per iteration (”expected” and ”original” in Fig. 12).
Figure 12. Iteration time of the LMT in Case 1.
@@
Problem 1: GPU throttling. PerfTracker outputs 300+ workers (located in 40+ hosts) with abnormal execution behaviors on multiple GPU computation kernel functions. Figure 1313(a) shows the example of GEMM (one of the GPU computation kernel functions). GPU computation kernels on these workers have much larger $\beta$ and smaller $\mu$ than most workers, so they have high values of $\Delta_{f,w}$. This indicates that the GPUs of these workers are much slower, so they need a longer time for computation. Interestingly, PerfTracker conducted multiple profiles of this LMT, and in each profile, the slow GPU workers were not consistent, but they are intensively distributed among workers in certain racks, rather than evenly distributed among all workers. This indicates that GPUs from specific deployment batches have intermittent performance issues. Hardware checks revealed recent GPU_NVSMI_HW_THROTTLE alerts on some of these machines, confirming the risk of intermittent GPU throttling. We have contacted the hardware vendor for repairs on these machines.
@@
Figure 13. Function execution behaviors of GEMM and AllGather_RING in Case 1. Each point represents the runtime behavior pattern of one worker. Outliers are regarded as abnormal function executions with high $\Delta_{f,w}$4.3) and are output by PerfTracker.
@@
Problem 2: NVLink down. PerfTracker also outputs 3 workers with abnormal execution behaviors on multiple collective communication functions. Taking AllGather_RING (one of the collective communication functions) as an example, 48 workers (belonging to three data-parallel groups, where each group has 16 workers) have significantly larger values of $\beta$ than the other 2,512 (Figure 13(b)). Among these 48 workers, three of them have significantly larger $\mu$ than the other 45 (Figure 1313(c)), indicating a higher bandwidth usage of PCIe. Hardware checks revealed NVLink Not Supported (NS) error on these three workers, which means all data flow sent to/from these three workers has to go through PCIe rather than NVLink. Since PCIe is much slower than NVLink, it leads to communication inefficiency.
@@
Fixes. We replaced the problematic hosts with standby hosts and restarted the LMT. Figure 12 shows the iteration time after the replacement (the ”fixed” line).
6.2. Case 2: Code-level Issues
A text-to-video LMT job of 3,072 Nvidia H800 GPUs is expected to take 3.5 seconds per iteration, but is training at 5 seconds per iteration (see the ”original” line in Figure 14).
Figure 14. Iteration time of the LMT in Case 2.
Problem 1: Low throughput in data loader. PerfTracker outputs abnormal function behaviors on the built-in function recv_into of the socket object (a low-level function call in the data loader) on multiple workers. As a GPU-independent function, we expect that the LMT is bottlenecked by it for no more than $1\%$ of the time ($\beta\leq 0.01$, as explained by “expected range” in §4.3). However, as shown in Figure 15(a), there are a lot of workers with a large value of $\beta$ for the recv_into function (so they have high value of $D_{f,w}$). This indicates low efficiency in data loading from the storage.
Problem 2: Inefficient implementation of Python function. Similar to the first problem, PerfTracker outputs abnormal behaviors on the Python function forward with large $\beta$ values, shown in Figure 15(b). This function performs CPU computation and launches GPU kernel functions. A large $\beta$ value indicates the function is bottlenecked by CPU computation, which needs to be optimized.
(a) recv_into of 3,072 workers.
Figure 15. CDF of $\beta$ values in two functions.
Problem 3: Asynchronous garbage collection. Apart from the above functions, PerfTracker also reports large $\beta$ values for some other Python functions like gradmode.py:__init__() and _flat_param.py: _get_unflat_views_unaligned. However, these functions are not CPU-intensive, and their executions with large $\beta$ are randomly distributed among all workers. This indicates these workers are executing Python garbage collection asynchronously. Each time a worker is collecting garbage, other workers have to wait for it, leading to GPUs idle and reduced training throughput.
Fixes. For the data loading problem, we suggest that the customer load the input data from our state-of-the-art parallel file system rather than the legacy object storage service, which can significantly reduce the time spent on data loading. For the Python garbage collection problem, we suggest that the customer manually call the garbage collection every 200 iterations. This guarantees all workers collect garbage at the same time, thus avoiding mutual waiting. Implementation optimization of the function forward is not trivial since it contains many CPU computation operations, and we are still working on it. At present, the iteration time of the LMT has reached $\sim$3.6 seconds, as shown in Figure 14.
6.3. System Overhead
Figure 16. PerfTracker’s overhead in timeline view.
Figure 16 illustrates PerfTracker’s overhead. When PerfTracker starts profiling, the LMT is executed with high-precision profiling for a 20-second window (§4.1). When the profiling window ends, LMT will be blocked by data generation for a short period. After that, LMT continues running with no more extra overhead, as the behavior pattern summary is done in a separate process (using a different CPU core), and the root cause localization is located remotely from the worker.
Overhead on production LMTs. We evaluate two LMTs: LMT-A of 2,560 GPUs in §6.1 and LMT-B of 3,072 GPUs in §6.2. Figure 17(a) shows the average iteration time during training with and without profiling. Figure 17(b) shows the duration of data generation, pattern summarization, and localization. The result shows that in production LMT, profiling does not affect the performance of LMT. After that, the localization can be finished within three minutes. In this process, the only overhead on LMT is data generation of $\sim$20 seconds, which is accepted by all of our customers.
(a) LMT iteration time with and without profiling.
(b) Time spent on different components.
(c) Time spent on localization under larger LMT scales.
Figure 17. The overhead introduced by PerfTracker.
Scalability to 1,000,000-GPUs LMTs. In PerfTracker, profiling, data generation and pattern summarization are all distributed in each worker’s container and executed in parallel, so their overhead is independent of the LMT scale. The only overhead proportional to the LMT scale is the centralized root cause localization, which is executed on a single CPU core. We generate simulated runtime behavior patterns as input, to evaluate the overhead of localization for larger-scale LMTs. Figure 17(c) shows that the localization requires only three minutes for a 1,000,000-GPU LMT. Together with the time spent on profiling, data generation, and pattern summarization, PerfTracker can make an end-to-end performance diagnosis for a 1,000,000-GPU LMT (which is larger than any existing LMT we have ever known) within 7 minutes. We cannot make a real evaluation of this, since currently we do not have a customer of such a training scale.

### v4

@@ 5 Implementation
We built Eroica with 7K lines of Python code, and deployed Eroica as a service for our production GPU clusters. We address the following practical challenges.
@@
Optimizations of profiling data generation. Eroica relies on Torch profiler to generate a part of raw profiling data (§4). However, Torch profiler introduces overhead to LMT: (1) When the profiling window ends, it requires a long time to prepare and dump the data, and (2) after profiling, it leaves a lot of resources (e.g., hooks for CUDA functions added by CUPTI [cupti]) in the LMT. They may reduce the performance of CUDA kernel execution even after profiling is finished.
@@
Eroica addresses the problems by optimizing the Torch profiler. First, the Torch profiler always transfers profiling data to Chrome tracing format, and then dumps the data using Kineto APIs [kineto]. Since Kineto supports dumping data with the same format, we remove the redundant and slow format transformation performed by Torch profiler, and directly apply Kineto to dump data. This reduces the time spent on data generation by 33%. Second, Eroica calls cuptiFinalize() to clear all remaining resources in LMT after profiling.
@@
Accessing hardware information within the user’s container. In production GPU clusters, an LMT typically runs inside containers; so does Eroica daemons (§4.1). However, GPU cluster providers usually restrict the permissions of user containers—Eroica cannot directly access hardware information. Eroica utilizes Kubernetes [kubernetes]’s native “emptyDir” feature to share directories between containers, performing high-frequency hardware sampling through a privileged management container to place data in the shared directory. Therefore, Eroica can obtain hardware profiling data within user containers without loosening permissions of user containers.
@@
Cooperating with other hardware monitors. In most production GPU clusters, coarse-grained hardware monitors are deployed on each physical host, using APIs such as DCGM and PCM to collect hardware metrics in real-time for health checks. However, some information (e.g., GPU metrics) can only be subscribed by one process at a time. Eroica’s data generator cooperates with the host’s monitoring system via signals (through a shared directory). As Eroica’s profiling typically lasts about 20 seconds each time, we do not encounter any conflicts caused by the cooperation.
6 Evaluation in Production
@@
Eroica has been deployed in our production GPU clusters with $\sim$100,000 GPUs for 1.5 years. During this period, there were in total 80 serious LMT performance issues for which we failed to localize the root causes using state-of-the-art techniques described in §2.2 (see Table 2). Eroica successfully identified root causes of 78 of them (success ratio=97.5%), improving the training throughput of these LMTs (the largest one includes 6,144 GPUs) by $\sim$20% to $\sim$100%.
@@
We present how we used Eroica to diagnose three LMTs with 3,072, 3,400, and 128 GPUs, respectively, to show:
@@
    Eroica can identify real-world performance issues of mixture of multiple hardware and code problems. (§6.1-6.2)
@@
    Eroica’s output can be directly fed into AI as the prompt to help customers automatically fix the code bug. (§6.3)
@@
    Eroica makes diagnosis in high efficiency, and does not introduce overhead to the routine training of LMT. (§6.4)
In addition, we present two more case studies in Appendices A and B, including one of the only two (out of 80) cases where Eroica failed to diagnose the issue. In Appendix C, we systematically compare Eroica with state-of-the-arts.
@@
Table 2: 80 serious performance issues identified by Eroica (ONLY includes those not identified by our existing systems)
Category Root cause Number of LMT cases
Hardware issues GPU 2
CPU 2
Network 6
Misconfigurations PyTorch 4
Communication 6
Dataloader 5
Low-efficiency code of users 45
6.1  Case Study 1: Code-level Issues
@@
A text-to-video LMT job on 3,072 Nvidia H800 GPUs is expected to take 3.5 seconds per iteration, but is training at 5 seconds per iteration (see the "original" line in Figure 12).
Figure 12: Iteration time of the LMT in Case Study 1.
Problem 1: Low socket throughput in data loader. Eroica outputs abnormal function behaviors on the built-in function recv_into of the socket object (a low-level function call in the data loader) on multiple workers. As a GPU-independent function, we expect that the LMT is bottlenecked by it in no more than $1\%$ of the time ($\beta\leq 0.01$, as explained by “expected range” in §4.3). However, as shown in Figure 13(a), there are a lot of workers with a high value of $\beta$ for the recv_into function (so they have a high value of $D_{f,w}$). This indicates low efficiency in data loading from the storage.
Problem 2: Inefficient implementation of Python function. Similar to the first problem, Eroica outputs abnormal behaviors on the Python function forward with large $\beta$ values, shown in Figure 13(b). This function performs CPU computation and launches GPU kernel functions. A large $\beta$ value indicates the function is bottlenecked by CPU computation, which needs to be optimized.
(a) recv_into of 3,072 workers.
Figure 13: CDF of $\beta$ values in two functions in Case Study 1.
Problem 3: Asynchronous garbage collection. Apart from the above functions, Eroica also reports large $\beta$ values for some other Python functions like gradmode.py:__init__() and _flat_param.py: _get_unflat_views_unaligned. However, these functions are not CPU-intensive, and their executions with large $\beta$ are randomly distributed among all workers. This indicates these workers are executing Python garbage collection asynchronously. Each time a worker is collecting garbage, other workers have to wait for it, leading to GPUs being idle and reduced training throughput.
Limitations of existing approaches. The customer did use Torch Profiler and Nsight System to profile their LMT job offline. As discussed in §2.1, profilers generate too much data; the common practice is profiling a few training iterations only in a fixed host (e.g., the rank-0 host). However, none of the three problems were detected, because in most cases the root cause of a low-performance user-level function is due to temporal pause on GPU/host memory allocation (e.g., garbage collection, GPU memory fragmentation, etc.) or access to remote storage. This only happens in a few random workers in each iteration, not all workers. Consequently, the offending function is difficult to detect without profiling all workers. In contrast, Eroica has a global view of function behaviors of all workers to identify abnormal ones with high efficiency.
Fixes. For the data loading problem, we helped the customer load the input data from our parallel file system rather than the legacy object storage service, which significantly accelerated data loading. For the Python garbage collection problem, the solution was to call the garbage collection every 200 iterations. This guaranteed all workers collect garbage at the same time, thus avoiding mutual waiting. Implementation optimization of the function forward is not trivial since it contains many CPU computation operations. At present, the iteration time of the LMT has reached $\sim$3.6 seconds, as shown in Figure 12.
6.2  Case Study 2: Mixed Code-Hardware Issues
A video generation LMT job on 3,400 Nvidia H800 GPUs is expected to take 8.5 seconds per iteration, but is training at 10.5 seconds per iteration ("expected" and "original" in Fig. 14). And the job crashes every few hours.
Figure 14: Iteration time of the LMT in Case Study 2.
(a) The $\beta$ value of SendRecv of 3,400 workers. This value is widely distributed across the 9%-16% range, with a small number (40 in total) of outliers distributed in 20%-23%.
(b) SendRecv’s $\beta$ and $\mu$ (GPU-NIC) of 40 workers with high $\beta$ (Outliers in Fig. 15(a)). Among the 40 workers, five have lower $\mu$ and one of them is significant lower which is the location of NIC down.
(c) The $\beta$ value of pin_memory of 3,400 workers. Most workers (3,397) is close to zero but three of them in range $23\%-33\%$.
(d) The $\beta$ and $\mu$ value of chunk_cat_cuda_kernel<float, c10::BFloat16> of 3,400 workers. All workers have similar $\mu$ but different $\beta$, indicating the workload imbalance.
Figure 15: Function execution behaviors in the Case Study 2.
Problem 1: Low cluster network throughput. As shown in Fig. 15(a), the $\beta$ value of most SendRecv functions ranges between 9% and 16%. However, our customer believes that the data amount in SendRecv is equal across all workers, so the $\beta$ values of SendRecv for all workers are expected to be identical. Moreover, based on the data amount in SendRecv and the NIC hardware, the $\beta$ value is expected to be $\sim$6%—indicating much lower cluster network throughput than expected. The root cause is that affinity-based flow scheduling is not deployed on this cluster, so inter-host data flow is not optimized.
Problem 2: NIC down. Fig. 15(a) shows another problem that the $\beta$ values of SendRecv of 40 workers (20%-23%) are much higher than most workers ($9\%-16\%$). Among the 40, one has high $\Delta_{f,w}$ due to the significantly lower $\mu$ than the other 39 (Fig. 15(b)), indicating a lower bandwidth between the GPU and the NIC. Our network engineers confirmed that the NIC corresponding to the worker is down.
Problem 3: Pin memory in ultra-high frequency. Fig. 15(c) shows that three random workers spent at most one third of time on pin_memory ($23\%-33\%$ $\beta$ value). Although this does not happen on the other 3,397 workers, they have to wait for the three workers to finish one iteration together, leading to reduced training throughput. Our customer confirmed that pin_memory operations come from data_loader processes.
Problem 4: Load imbalance. Eroica finds that almost all GPU computation kernel functions have similar $\mu$ but significantly different $\beta$ across workers. Fig. 15(d) shows one GPU kernel function for example. The busiest GPU spends 46% more time computing than the most idle one. This indicates that although workers execute GPU kernels with equal performance, some workers launch much more GPU kernels than others, leading to load imbalance. Our customer confirmed that the model is a video-to-video model, for each training iteration, video inputs for workers have different duration, requiring different amount of GPU kernel launches. Workers with less GPU computation have to wait for others with more GPU computation to finish one iteration together; this leads to poor training efficiency (e.g., low Model FLOPs Utilization).
Limitations of existing approaches. For Problem 1, hardware monitoring on the cluster network showed no warning. The problem was low-efficiency flow scheduling that failed to fully utilize the network. For Problem 2, the affected host was newly added to the cluster, and the monitoring agent had not been updated, therefore, failed to alarm the problem. For Problem 3, the pure code-level problem, it only happened on three of 3,400 workers in one iteration, so it was hard to be captured offline (as explained in §6.1). For Problem 4, although a GPU may receive a long video in one iteration to reach high utilization, due to the randomness of input scheduling, it does not happen in all iterations. Based on our GPU utilization monitor, in general all GPUs have similar utilization in a period of time, so it failed to detect the problem.
Fixes. We removed 20 hosts with the least network throughput in Fig. 15(a) to mitigate Problem 1, including the host with NIC down (Problem 2), then the iteration time reduced from 10.5s to 9.5s ("hw_fix" line in Fig. 14). For Problem 3, our customer decreased the number of data_loader processes to reduce the memory overload. It not only decreased the iteration time (from 9.5s to 9.2s), but also solved the crash problem. For Problem 4, we helped the customer balance the input between workers. Currently, this LMT is training at 8.5 seconds per iteration (the "all_fixed" line in Fig. 14) without crashes. With all of the above efforts, the input data processing rate was improved from 7,173 samples/day to 9,644 samples/day, a 34% end-to-end performance improvement.
6.3  Case Study 3: Diagnose and auto-fix customer-code bugs with AI support
A robotics (embodied AI) model’s training of 128 GPUs got stuck for a long time, and the customer requested a diagnosis.
Problem detected: Stucked in dataset preloading. In this stalled training job, Eroica found a single worker behaving differently from the rest. On this worker, a data-loading/preprocessing thread is blocked in queue.put() inside dynamic_robot_dataset._preload() (waiting on a Python thread lock), indicating the input pipeline is stuck/back-pressured. Meanwhile, other workers are either sleeping in dataset management routines (e.g., _monitor_config, _run_threads) or waiting in JAX [jax] execution, suggesting they are idle and effectively waiting for this worker to make progress. Overall, the hang is most consistent with a Python-side data pipeline deadlock or queue blockage in the dataset prefetch/preload logic.
Automatically fixed by AI. We instructed the customer to use Eroica’s output (the stuck Python function) together with the relevant code as a prompt to Cursor [cursor] (Sonnet 4.5), an AI coding tool. Cursor immediately identified the bug: during data fetching, a logging/debug print accessed array[0] on a sharded distributed array, implicitly triggering an all-gather. Since this collective was invoked at an unexpected point (not executed by all ranks), it led to a distributed deadlock and stalled the training job. Cursor then automatically generated the fix for the buggy code, and the training job resumed normally.
6.4  System Overhead
Figure 16: Eroica’s overhead in timeline view.
Figure 16 illustrates Eroica’s overhead. When Eroica starts profiling, the LMT is executed with high-precision profiling for a 20-second window (§4.1). When the profiling window ends, LMT will be blocked by data generation for a short period. After that, LMT continues running with no extra overhead, as the behavior pattern summarization is done in a separate process (using a different CPU core), and the root cause localization is performed remotely from the worker.
Overhead on production LMTs. We evaluate two LMTs: LMT-A of 3,072 GPUs in §6.1 and LMT-B of 3,400 GPUs in §6.2. Figure 17(a) shows the average iteration time during training, with and without profiling. Figure 17(b) shows the duration of data generation, pattern summarization, and localization. The results show that profiling does not affect the performance of production LMTs; summarization and localization complete within three minutes. The main overhead on LMT is data generation ($\sim$20) seconds, which is incurred just once and accepted by all our customers. Notably, pattern summarization, uploading, and root cause localization occur outside the training process. Although these steps may last for minutes, they introduce zero overhead to the training task. Regarding memory usage, although profiling consumes tens of GBs on each host, this is negligible given that production training hosts typically have 1–2 TB of host memory.
(a) LMT iteration time with and without profiling.
(b) Time spent on different components.
(c) Time spent on localization under larger LMT scales.
Figure 17: The overhead introduced by Eroica.
We also evaluate the overhead with different models and configurations in Appendix D to show Eroica is able to make diagnosis with high efficiency in diverse situations.
Scalability to 1,000,000-GPUs LMTs. In Eroica, profiling, data generation and pattern summarization are all distributed in each worker’s container and executed in parallel, so their overhead is independent of the LMT scale. The only overhead proportional to the LMT scale is root-cause localization, which is executed on a single CPU core. We generate simulated runtime behavior patterns as input, to evaluate the overhead of localization for larger-scale LMTs. Figure 17(c) shows that the localization requires only three minutes for a 1,000,000-GPU LMT. Together with the time spent on profiling, data generation, and pattern summarization, Eroica can make an end-to-end performance analysis for a 1,000,000-GPU LMT within 7 minutes.
7 Discussions: Enhancing Eroica with AIOps
Eroica does not guarantee automatic identification of the root causes for all possible performance issues. (1) For most performance issues, Eroica’s output of the problematic functions and their corresponding runtime behavior patterns is sufficient to identify the root cause or can be fed into AI to determine the root cause (like the case in §6.3). However, some user-defined Python functions and CPU kernels have complex code and logic. Therefore, even when Eroica identifies that these functions have problems (e.g., excessively long execution times), analyzing the root causes may still require manual inspection of their code. (2) Eroica does not ensure detecting problems outside the training task. For example, a background process running on the physical host may incidentally consume too many resources, leading to a performance issue for the training task in the container on that host. A case study is presented in Appendix B.
To bridge the last-mile gap between abnormal function behavior and the root cause, we are instructing our customers to combine Eroica’s output with additional information (the code of the function with abnormal behavior, background processes with system call activities, and hardware configuration/utilization) and generate a standardized prompt for an AI model to diagnose the root cause. In production, while the AI provides correct diagnoses only in a subset of cases, it yields useful hints in most cases, which is particularly valuable for performance issues caused by user code. Therefore, in our research on AIOps (Artificial Intelligence for IT operations) for LMT, Eroica’s output serves as the core component for prompt construction.
8 Related Work
Hardware monitoring tools. Accelerators expose rich telemetry through vendor tools. For NVIDIA GPUs, NVML [nvml] provides a C-based API to query and sample device status (e.g., utilization, temperature, power, memory, and error states). DCGM [dcgm] targets large clusters by offering centralized, scalable GPU health/performance monitoring and diagnostics for fleet management. CPU performance and energy metrics can be collected via Intel PCM [pcm] (e.g., cache and memory bandwidth statistics). For NICs (e.g., Mellanox), mstflint [mstflint] provides access to device counters such as packet/byte throughput and link status.
LMT profiling and analysis. Nsight Systems and Nsight Compute provide complementary profiling for CUDA workloads. Nsight Systems captures an end-to-end timeline across CPU threads, GPU work, and memory transfers, whereas Nsight Compute focuses on per-kernel metrics (e.g., occupancy and memory throughput). Torch Profiler, integrated into PyTorch, reports fine-grained per-operator/per-function events across Python, CPU, and CUDA to attribute training time and identify bottlenecks.
9 Conclusion
This paper presents Eroica, an online performance troubleshooting system for LMT. Eroica presents the process of profiling on performance degradation, behavior pattern summarization, and performance issue localization, which utilize runtime behavior patterns of functions to make diagnosis. We deploy Eroica in real-world production GPU clusters and solve sophisticated performance issues for real users.
Acknowledgements
We thank our shepherd Changhoon Kim, and the anonymous reviewers for their insightful comments. Ennan Zhai is the corresponding author.
APPENDIX
Appendix A Case Study 4: Hardware Issues
A text-to-picture LMT job of 2,560 Nvidia H800 GPUs is expected to take 5 seconds per iteration, but is training at 9 seconds per iteration ("expected" and "original" in Fig. 18).
Figure 18: Iteration time of the LMT in Case Study 4.
@@ Appendix A Case Study 4: Hardware Issues
Problem 1: GPU throttling. Eroica outputs 300+ workers (located in 40+ hosts) with abnormal execution behaviors on multiple GPU computation kernel functions. Figure 1919(a) shows the example of GEMM (one of the GPU computation kernel functions). GPU computation kernels on these workers have much larger $\beta$ and smaller $\mu$ than most workers, so they have high values of $\Delta_{f,w}$. This indicates that the GPUs of these workers are much slower, so they need a longer time for computation. Interestingly, Eroica conducted multiple profiles of this LMT, and in each profile, the slow GPU workers were not consistent, but they are intensively distributed among workers in certain racks, rather than evenly distributed among all workers. This indicates that GPUs from specific deployment batches have intermittent performance issues. Hardware checks revealed recent GPU_NVSMI_HW_THROTTLE alerts on some of these machines, confirming the risk of intermittent GPU throttling. We have contacted the hardware vendor for repairs on these machines.
@@
Figure 19: Function execution behaviors of GEMM and AllGather_RING in Case Study 3. Each point represents the runtime behavior pattern of one worker. Outliers are regarded as abnormal function executions with high $\Delta_{f,w}$4.3) and are output by Eroica.
@@
Problem 2: NVLink down. Eroica also outputs 3 workers with abnormal execution behaviors on multiple collective communication functions. Taking AllGather_RING (one of the collective communication functions) as an example, 48 workers (belonging to three data-parallel groups, where each group has 16 workers) have significantly larger values of $\beta$ than the other 2,512 (Figure 19(b)). Among these 48 workers, three of them have significantly larger $\mu$ than the other 45 (Figure 1919(c)), indicating a higher bandwidth usage of PCIe. Hardware checks revealed NVLink Not Supported (NS) error on these three workers, which means all data flow sent to/from these three workers has to go through PCIe rather than NVLink. Since PCIe is much slower than NVLink, it leads to communication inefficiency.
Limitations of hardware monitors. Section 2.1 discusses limitations of hardware monitoring. In this case, for Problem 1, the affected GPU throttling events occur at low frequency with short durations (100us–10ms), failing to trigger Nvidia’s built-in alerts. The monitoring system, operates at minute-level granularity (with second-level precision in short intervals) to be light-weighted, also failed to capture such fine-grained performance fluctuations. For Problem 2, the affected machines were newly added to the cluster, and the monitoring agent had not been updated, thus failing to detect the problem.
@@
Fixes. We replaced the problematic hosts with standby hosts and restarted the LMT. Figure 18 shows the iteration time after the replacement (the "fixed" line).
Appendix B Case Study 5: Code Issue (Eroica Failed to diagnose)
A reinforcement learning LMT job of 8 Nvidia H800 GPUs (located at one host) is expected to take $\sim 22$ seconds per training iteration (about hundreds of commits before the current version, called Version A), but is training at $\sim 26$ seconds per training iteration (called version B).
(a) $\beta$ value comparison of three representative GPU computation kernel functions between Version A and B.
(b) $\beta$ value comparison of representative collective communication functions between Version A and B.
Figure 20: $\beta$ values of GPU computation and communication functions in Case Study 4 (we omit kernel names of GPU computation kernel functions since they are too long).
Figure 21: The runtime behavior of an MoE training.
Figure 22: The python function behaviors of a forward step in the MoE training.
Figure 23: The kernel function (includes both computation and communication) behaviors of a forward step in the MoE training.
Eroica’s Diagnosis result. Eroica diagnosed both Version A and B, revealing that most GPU computation kernels and collective communications in Version B exhibited slightly higher $\beta$ values than Version A (Fig. 20(a) and 20(b)). Since $\beta$ represents the time proportion within a single iteration and Version B had longer iteration durations, this indicates that most computation and communication functions consumed more time in Version B. In addition, Eroica observed no significant difference in $\mu$ values between the two versions, which confirmed no hardware issues, indicating that Version B executed more workload on GPUs and networks during identical functions.
However, our customer confirmed that the workload of Version A and B were exactly the same.
The root cause. This LMT job utilizes reinforcement learning techniques to train the data. It has both training processes and inference processes. Inference processes should execute remotely but also persist in an idle state on the host due to developer oversight, running concurrently with training processes while repeatedly invoking allgather for synchronization. Inference processes utilize gloo [gloo] to execute allgather by default, which is based on TCP, so they do not influence the training performance. However, in one commit (among the hundreds of commits between Version A and B), it was changed to NCCL (based on CUDA) to pursue better performance. The side-effect is that it contended for GPU resources with training processes (NCCL requires GPU SMs). This led to lower performance of both GPU computation and communication of training processes.
Why Eroica failed and how the root cause is identified. As discussed in §7, for most performance issues, when the problematic function and its behavior are identified, it is very close to the root cause, or we can even directly derive the root cause. In this case, however, there were too many "problematic" functions (including most GPU computation and communication functions). Moreover, no one was aware that training and inference are executed in the same host (in separate processes) in accident. Since only training processes have performance issues, we overlooked employing Eroica to diagnose other processes (e.g., the idle inference process) on the same host.
The root cause is finally identified by human power. A team of 20 engineers performed a binary search among hundreds of commits between Version A and B, identifying a performance difference between two adjacent commits. This localization process took one month.
Lessons and rethinking: Potential opportunities to automatically diagnose this case. As reinforcement learning becomes a promising technique of LMT, some LMT jobs execute processes of multiple actors in one host, like training and inference.
Based on the identified root cause, we retrospectively analyzed the diagnostic process. Eroica had at least two opportunities to successfully locate this issue: (1) We should first make an overview at all processes running on the host. Upon recognizing the potential concurrency between the (idle) inference and training processes in this LMT, Eroica should have been deployed to diagnose the idle inference process also, which would have directly revealed the NCCL function calls in the inference process. (2) Even based solely on training process diagnostics, the root cause could have been indirectly deduced. Eroica’s detection of heavier workload execution in Version B versus Version A functions indicated resource contention.
The failure in problem localization primarily originated from our lack of an overview of the LMT, causing us to overlook the impact of other processes (even existing accidentally) on the training procedure. For future performance diagnostics, especially for the LMT with complex inter-process relationships, Eroica will implement usability optimizations, such as automatically expanding the diagnosis scope to all LMT-related processes within the host.
Appendix C Comparisons with State of the Arts
We compare Eroica with recent monitoring techniques [jiang2024megascale, ncclprofiler, bpftrace] and profilers; Table 3 summarizes the results.
For online monitors, MegaScale [jiang2024megascale] deploys techniques to record CUDA-event timelines (to expose slow GPU kernels) and perform ms–s RDMA monitoring, but mainly report alerts/statistics; root-cause localization (especially for network issues) is typically manual. It also lacks Python function events, and thus cannot diagnose code-level issues. NCCL Profiler targets only communication by instrumenting the communication library. Bpftrace [bpftrace] selectively instruments key Python functions/kernels (e.g., via .so replacement) to detect slowdowns, but root-cause analysis still often requires manual effort. We also compare against directly using Nsight Systems [nsys] and Torch Profiler [torch_profiler]. Although they support offline/custom analysis (Nsight’s Post-Collection Analysis API and Torch’s JSON dumps), running them online incurs prohibitive overhead in production LMT.
Table 3 further compares troubleshooting ability and diagnosis time on the problems in §6.1 and §6.2 (e.g., Case1-P1 represents Problem 1 of Case Study 1). Existing online monitors miss many issues due to incomplete data sources. Profiler traces cover more, but loading data from all workers can take days; since they do not provide built-in troubleshooting logic, we report only the data-loading time (end-to-end diagnosis would take longer).
Table 3: Comparisons with state-of-the-art approaches
Technique Case 1 Case 2 Diagnostic time for 10,000-GPU LMT
P1 P2 P3 P1 P2 P3 P4
MegaScale [jiang2024megascale] online
NCCL Profiler [ncclprofiler] online
bpftrace [bpftrace] online
Nsight System >1.5 days (offline)
Torch Profiler >3.5 days (offline)
Eroica 3 min (online)
Appendix D Profiling Overhead in Different Model Configurations